If the width of a rectangle is 5 meters less than the length and the perimeter of the rectangle is 53 meters, find the length and width. could you show me how to do this?
Answer by stanbon(75887)

If the width of a rectangle is 5 meters less than the length and the perimeter of the rectangle is 53 meters, find the length and width.
---------------
Let length be "x".
Then width is "x-5"
----
Perimeter = 2(length + width)
-----
53 = 2(x + x-5)
53 = 2(2x-5)
53 = 4x-10
4x = 63
x = 15.75 meters (length)
x-5 = 10.75 meters (width)
==============================
Cheers,
Stan H.
===============